OpenStudy (anonymous):

find the area of a rectangular object which has the length of 5 √8 feet and a with of 2 √ 5 feet

OpenStudy (campbell_st):

well area is A = length x width so you are multiplying radicals \[A = 5\sqrt{8} \times 2\sqrt{5} = 5 \times 2 \times \sqrt{8} \times \sqrt{5} \] all you need to do is simplify the product of the radicals.

OpenStudy (anonymous):

how exactly do i do that?

OpenStudy (campbell_st):

use one of the rules for radicals. \[\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\]

OpenStudy (anonymous):

im sorry but i still dont have a clue

OpenStudy (campbell_st):

ok... so \[\sqrt{8} \times \sqrt{5} = \sqrt{8 \times 5} \] so the area is \[A = 5 \times 2 \times \sqrt{8 \times 5}\] hope this helps.

OpenStudy (anonymous):

yes it does thanks a bunch